OpenStudy (anonymous):

How would one solve x^2+5=4x?

OpenStudy (akashdeepdeb):

x^2 - 4x + 5 = 0 [a=1, b = -4, c = 5] Use the quadratic equation: \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] Got it? :)

OpenStudy (anonymous):

Ah, that makes more sense. Thank you very much!

OpenStudy (akashdeepdeb):

:)